OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2 i got y = 2x + 5

OpenStudy (anonymous):

eq of line in slope-point form is given as y-y1=m(x-x1) where m=slope= 2 x1= 3 and y1=-2 Thus y-y1=m(x-x1) reduces to y-(-2)=2(x-3) i.e. y+2=2x-6 i.e. y= 2x-6-2 i.e. y= 2x-8 which is in the slope -intercept form